Determine the current flowing through an element if the charge flow is given by a. q(t)= (3e-t -5e-2t) nC b. q(t)= 10sin120πt pC c. q(t)= 20e-4tcos50t µC

1
Expert's answer
2021-11-01T17:11:19-0400

If the charge flow is given as a function of time, i.e. q(t), we determine the current flowing through an element by:

"I = \frac {dq(t)} {dt}"

where I = current


(a) For the charge flow given by: q(t)= (3e-t -5e-2t) nC;

"I = \frac {dq(t)} {dt} \\
I = \frac {d(3e-t -5e-2t) nC} {dt} \\
I = \frac {d} {dt} (3e-t -5e-2t) nC \\"

differentiating q(t) = (3e-t -5e-2t) nC with respect to time (t),

"I = \frac {d} {dt} (3e) - \frac {d} {dt} (t) - \frac {d} {dt} (5e)- \frac {d} {dt} (2t) \ nC \\
I = (0 - 1 - 0 - 2) nC \\ 
I = -3 nC \\
or \ I = -3 \times 10^{-9} \ C"


(b) For the charge flow given by: q(t) = 10sin120πt pC

"I = \frac {dq(t)} {dt} \\
I = \frac {d(10 \sin120\pi t) \ pC} {dt}"

differentiating q(t) = 10sin120πt pC with respect to time (t),

"I = \frac {d} {dt} (10 \sin 120\pi t) pC \\
let \ w = 120\pi t \\ u= \sin w \\ and, \ q = 10u \\
I = \frac {dq} {dt} = \frac {dw} {dt} \times \frac {du} {dw} \times \frac {dq} {du} \\
w = 120 \pi t, \ \frac {dw} {dt} = \frac {d} {dt} (120\pi t) = 120 \pi \\
u = \sin w, \frac{du}{dw} = \frac {d} {dw} (\sin w) = \cos w \\
q = 10 u, \frac {dq} {du} = \frac {d} {du} (10u) = 10 \\
I = \frac {dq} {dt} = 120 \pi \times \cos w \times 10 \ pC \\
I = 1200 \pi \cos 120 \pi t \ pC \\
or \\ 
I = 1200 \pi \cos 120 \pi t \times 10^{-12} \ C \\
I = 1.2 \pi \cos 120 \pi t \times 10^{-9} \ C \\
I = 1.2 \pi \cos 120 \pi t \ nC"


(c) For the charge flow given by: q(t)= 20e-4tcos50t µC

"I = \frac {dq(t)} {dt} \\
I = \frac {d(20e-4tcos50t) \ µC} {dt} \\"

differentiating q(t) = 20e-4tcos50t µC with respect to time (t),

"I = \frac {d} {dt} (20e-4t\cos50t) \ µC \\
I = \frac {d} {dt} (20e) - \frac {d} {dt} 4t\cos50t \\
\frac {d} {dt} (20e) = 0 \\
\therefore I = - \frac {d} {dt} (4t\cos50t) \\
let \ w = 50 t \\ u= \cos w \\ and, \ q = 4tu \\
I = \frac {dq} {dt} = \frac {dw} {dt} \times \frac {du} {dw} \times \frac {dq} {du} \\
w = 50 t, \ \frac {dw} {dt} = \frac {d} {dt} (50 t) = 50 \\
u = \cos w, \frac{du}{dw} = \frac {d} {dw} (\cos w) = -\sin w \\
q = 4tu, \frac {dq} {du} = \frac {d} {du} (4tu) = 4t \\
I = 0 - \frac {dq} {dt} = - \frac {dq} {dt} = - (50 \times -\sin w \times 4t) \ µC \\
I = 200t \sin 50 t \ µC \\
or \\ 
I = 200t \sin 50 t \times 10^{-6} \ C \\
I = 0.2t \sin 50 t \times 10^{-3} \ C \\
I = 0.2t \cos 50 t \ mC"